conservation of energy chapter

A 62-kg skier starts from rest at the top of a ski jump, point A inFig. 8-41, and travels down the ramp.If friction and air resistance can be neglected, 

(a)determine her speed vwhen she reaches the horizontal end of the ramp at B.

(b)determine the distance s to where she strikes the ground at C.
Answer #1
A 62-kg skier starts from rest at the top of a ski jump, point A inFig. 8-41, and travels down the ramp. If friction and airresistance can beneglected, (a) determine her speed vbwhen she reaches the horizontal end of the ramp at B.

v=√0+2*9.8*40.6=28.20m/sec.

(b) determine the distance s to where she strikes the groundat C.
when it covers 4.4m vertical height at the end of B the speed=√28.20*28.20+2*9.8*4.4=29.68m/sec
at the end of B vertical component ofvelocity=29.68sin30=14.84m/sec.
now let the vertical distance to be covered be h
now time =h=14.84t+4.9t*t
calculate time from this equation
after that horizontal distance=29.68cos30*t which is thedesired answer.





answered by: My Help Pls1